OpenStudy (anonymous):

Find the limit of each function; a) as x approaches infinity b) as x approaches negative infinity f(x)=(2/x)-3

OpenStudy (anonymous):

both a and b = -3

OpenStudy (anonymous):

@Hoa I am confused how both equals -3?

OpenStudy (johnweldon1993):

is that \[(2/x) - 3\] or is that "-3" an exponent?

OpenStudy (johnweldon1993):

I'll assume it is the first one.....(2/x) - 3 now ask yourself....what happens as x gets infinitely big? well since its in the denominator....the number gets closer and closer to 0 right? think about 1/10....thats .10....now 1/100...thats .01....1/1000....that .01 ...closer and closer to 0 right? so if 2/x...gets closer and closer to zero.....and the 3 is subtracted from it......that would approach -3 right? now with - infinity....the same thing......we still approach 0...just from a different side......then still subtract 3....so we still approach -3

OpenStudy (anonymous):

lim 2/x \[\lim_{x \rightarrow \infty}(\frac{ 2 }{ x } -3 ) = \lim_{x \rightarrow \infty}\frac{ 2 }{ x } -\lim_{x \rightarrow \infty}3\]= 0 -3 = -3

OpenStudy (anonymous):

the same thing with - infinitive.

OpenStudy (anonymous):

is it your problem or else? (2/x) -3 or (2/x)^-3? if it's latter, the things is different.

OpenStudy (anonymous):

if yours is \[(\frac{ 2 }{ x })^-3\] then the answer is infinitive or - infinitive respectively

OpenStudy (anonymous):

do you know why? because (2/x)^-3 = x^3 /2. x--> infinitive , limit of x is approaches infinitive, too.

OpenStudy (anonymous):

sorry, x^3/8

OpenStudy (anonymous):

@johnweldon1993 its like this in the input 

OpenStudy (johnweldon1993):

yeah exactly...that's what I assumed....then my explanation stands :)